OpenStudy (anonymous):

What value of c solves the equation? 4^c=1/16 c =

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

What this question is asking is, 4 raised to what power will equal 1/16.

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

4^c=(2^2)^c=2^(2c)=1/16=2^-4 2c=-4 c=-2